# Supplement

### v3

@@ Abstract
Functional Principal Components Analysis (FPCA) is a widely used analytic tool for dimension reduction of functional data. Traditional implementations of FPCA estimate the principal components from the data, then treat these estimates as fixed in subsequent analyses. To account for the uncertainty of PC estimates, we propose FAST, a fully-Bayesian FPCA with three core components: (1) projection of eigenfunctions onto an orthonormal spline basis; (2) efficient sampling of the orthonormal spline coefficient matrix using polar decomposition; and (3) ordering eigenvalues during sampling. Extensive simulation studies show that FAST is very stable and performs better compared to existing methods. FAST is motivated by and applied to a study of the variability in mealtime glucose from the Dietary Approaches to Stop Hypertension for Diabetes Continuous Glucose Monitoring (DASH4D CGM) study. All relevant STAN code and simulation routines are available as supplementary material.
@@ 1 Introduction
Functional principal component analysis (FPCA) is a common analytic approach for high-dimensional applications. FPCA approximates the covariance between all possible pairs of observations by identifying a small set of orthonormal principal components (FPCs). Using FPCs to provide a reduced-rank representation of functional data, rather than fixed (spline) bases (Shi et al.,, 1996; Rice and Wu,, 2001), provides a data-driven representation based the leading modes of variability (James et al.,, 2000). While the means of estimating these FPCs can vary, e.g. polynomial smoothing (Yao et al.,, 2005), kernel smoothing (Staniswalis et al.,, 1998), or sandwich covariance smoothing (Xiao et al.,, 2016), the key idea remains consistent. Conditioning on the FPCs, functional models become mixed effects models with a small number of random effects and inference can be conducted using existing software (Reiss and Ogden,, 2007; Liu and Guo,, 2012; Di et al.,, 2009; Zhou et al.,, 2010; Shou et al.,, 2015). Frequentist implementations of FPCA follow a two-step approach: (1) diagonalizing a smooth estimator of the observed functional covariance; and then (2) conducting inference conditional on the FPC estimates (Xiao et al.,, 2016; Crainiceanu and Goldsmith,, 2010; Peng and Paul,, 2009). This conditional approach may lead to invalid inference in smaller samples and/or when eigenfunctions explain smaller variance (Goldsmith et al.,, 2013). FPC forms may be falsely identified, or FPC scores with substantial uncertainty may be used in downstream regression.
@@
There are several Bayesian FPCA implementations which jointly model the FPCs and their scores to account for all sources of uncertainty. An early fully-Bayesian FPCA was developed by Goldsmith et al., (2015), who modeled both the fixed and random effects functions using unconstrained splines and applied post-processing to obtain orthonormal FPC samples. Jauch et al., (2021) directly sampled the FPC matrix using polar decomposition and Gaussian Process priors. In the context of sparse functional data, Ye, (2024) projected the FPCs onto an orthonormal spline basis and used the Gibbs sampling approach of Hoff, (2009) to sample the spline weights such that FPC orthonormality was preserved. A recent approach used variational inference based on message passing (Nolan et al.,, 2023).
@@
In this paper, we introduce Fast Bayesian FPCA, or FAST for short, which integrates features of Ye, (2024) and Jauch et al., (2021) along with computational optimizations to produce an MCMC approach that is stable, scalable, and extensible. FAST leverages orthogonal spline expansion to reduce the dimension of sampled FPC parameters to the corresponding weights, which are then efficiently sampled using polar decomposition. The dimensionality reduction of the posterior space by spline expansion combined with stabilizing order constraints of FPCA ensures that the approach has excellent convergence and mixing properties while being computationally efficient. All core components of FAST are easy to implement in standard Bayesian modeling software such as STAN (Carpenter et al.,, 2017) and can be extended to include covariates, multilevel, sparse, and multivariate data.
@@ 2.1 FPCA model
The functional data structure has the form $Y_{i}(t),i=1,\ldots,N$ for $t\in\{t_{1},\ldots,t_{M}\}\in[0,1]$, which can be thought of as latent functions observed on a finite grid. Assuming that $Y_{i}(t)$ are generated by the same Gaussian Process and truncating the decomposition (Karhunen,, 1947; Kosambi,, 1943; Loève,, 1978) produces the standard FPCA model:
@@
Model (1) is a first line approach to analysis of high dimensional data (Crainiceanu et al.,, 2024) because in many applications $K\ll M$ and model (1) provides a practical, low-dimensional approximation to the original high-dimensional data. Moreover, conditional on the FPCs, $\phi_{k}(\cdot)$, the model becomes a mixed effects model with uncorrelated random effects. In this paper we treat the eigenfunctions $\phi_{k}(t)$ as parameters, estimate them from the data, and account for uncertainty. FAST provides a stable approach to simulating the posterior distribution of all parameters, including the eigenfunctions.
@@ 2.2 Orthonormal spline expansion
[…]
@@
We will show that the $\phi_{k}(t)$ functions are orthonormal if and only if $\boldsymbol{\Psi}^{t}\boldsymbol{\Psi}=I_{K\times K}$, where $\mathbf{\Psi}=[\psi_{1}|\ldots,|\psi_{K}]$ is the $Q\times K$ dimensional matrix obtained by column binding the $\psi_{k}$. Indeed, as $\mathbf{B}(t)$ is an orthonormal spline basis, $\langle B_{i}(t),B_{j}(t)\rangle=1$ if $i=j$ and 00 otherwise. Therefore, $\langle\phi_{k}(t),\phi_{k^{\prime}}(t)\rangle$ is equal to
[…]
@@
The quantity $\psi_{k}^{t}\psi_{k^{\prime}}=1$ when $k=k^{\prime}$ and 00 if and only if $\boldsymbol{\Psi}$ matrix is orthonormal. Therefore, the eigenfunctions $\phi_{k}(t)$ are orthonormal in $L^{2}([0,1])$ if and only if the low-dimensional spline coefficients are orthonormal in $\mathbb{R}^{Q}$. This is a crucial point of our methodology, as the orthonormality of the eigenfunctions (in infinite or high dimensional spaces) is controlled by orthonormality of vectors in a small dimensional space. Because it is much easier the sample from a small dimensional space, this projection is one of the main reasons the FAST method is stable and computationally efficient.
@@
In functional data analysis, orthonormality in the theoretical $L^{2}([0,1])$ functional space is often confounded with orthonormality in the $\mathbb{R}^{M}$ observed vector space. For a fixed $M$, the difference is minimal, as this corresponds to approximating the functional inner product with the corresponding Riemann sum – the inner product in $\mathbb{R}^{M}$ scaled by $1/M$. However, when the dimension is changed from $M_{1}$ to $M_{2}$, the vector inner product must be adjusted by $\sqrt{M_{1}/M_{2}}$, whereas the functional inner product is not affected. FAST can be used in the vector space as well by choosing matrix $\mathbf{B}$ such that $\mathbf{B}^{t}\mathbf{B}=\mathbf{I}_{Q}$. The resulting FPC matrix $\mathbf{B}\boldsymbol{\Psi}$ is then a product of orthonormal matrices. To maintain robustness to different $M$, we focus here on the functional orthonormality. However, the software implementation in supplementary materials can leverage vector orthormality, as well.
@@
The basis dimension $Q$ must be sufficiently large to capture the maximum complexity of the subject-specific functions. We follow here the standard recommendations for penalized splines and use $Q\in[20,40]$ (Ruppert,, 2002), which performed well in our tests (see Supplement Sections S4.2). The choice of the basis matrix $\mathbf{B}$ can substantially impact the computational stability and efficiency of FAST. While, in theory, any spline matrix $\mathbf{B}$ which is orthonormal in $L^{2}[0,1]$ could be used, in practice some approaches have better computational properties and interpretation properties. We choose as our basis ‘Splinets’, an orthonormalized B-spline basis designed by Liu et al., (2020) to maintain temporal localization. We augment their basis with a slope and intercept, appropriately orthonormalized.
@@ 2.3 Priors
[…]
$f(w_{\mu})\propto h_{\mu}^{\text{R}(\mathbf{P}_{\alpha})/2}\exp\left\{-\frac{h_{\mu}}{2}w^{t}_{\mu}\mathbf{P}_{\alpha}w_{\mu}\right\}\;,\;\;f(\psi_{k})\propto h_{k}^{\text{R}(\mathbf{P}_{\alpha})/2}\exp\left\{-\frac{h_{k}}{2}\psi^{t}_{k}\mathbf{P}_{\alpha}\psi_{k}\right\}\;,$
where rank $\text{R}(\mathbf{P}_{\alpha})$ is the rank of the matrix $\mathbf{P}_{\alpha}$ and $h_{\mu},h_{k}>0$ are the smoothing parameters for $\mu(\cdot)$ and $\phi_{(}\cdot)$, respectively. FAST uses separate parameters for each functional component, in-line with Ye, (2024) and Goldsmith et al., (2015).
We assume that the inverse variance components, including the smoothing parameters, have non-informative Gamma priors with the shape and rate equal to $0.001$. For discussion on the choice of Gamma priors for inverse variance components, see Crainiceanu et al., (2005); Crainiceanu and Goldsmith, (2010)
@@ 2.4 Polar decomposition
With these priors, the conditional posterior distributions for all parameters except $\boldsymbol{\Psi}$ have analytical forms; see Supplement Section S2. The $Q\times K$ dimensional matrix parameter $\boldsymbol{\Psi}$ is an element in the space of orthonormal matrices. This space is called the Stiefel manifold (James,, 1976), has a finite volume, and is denoted as $\mathcal{V}_{K,Q}$ (Chikuse,, 2003). We assume a uniform prior for $\boldsymbol{\Psi}$ over $\mathcal{V}_{K,Q}$, with density $\mathbbm{1}(\mathbf{\Psi}\in\mathcal{V}_{K,Q})$ (Mardia and Khatri,, 1977).
Result 1.
@@
The distribution in Result 1 does not follow a known form, though it is close to the generalized Langevin-Bingham family (Hoff,, 2009). We sample this distribution indirectly using the polar decomposition of matrices. The polar decomposition of the $Q\times K$ dimensional matrix $\mathbf{X}$ is $\mathbf{X}=\mathbf{UP}$, where $\mathbf{U}\in\mathbb{R}^{Q\times K}$ is orthonormal (on the Stiefel manifold $\mathcal{V}_{K,Q}$) and $\mathbf{P}\in\mathbb{R}^{K\times K}$ is positive semi-definite. The polar decomposition is unique when $\mathbf{X}$ is full rank $K$. Crucially, if $\mathbf{X}$ has independent $N(0,1)$ entries, then the matrix $\mathbf{U}$ is uniformly distributed over the Stiefel manifold $\mathcal{V}_{K,Q}$ (Chikuse,, 2003). Therefore, the matrix $\boldsymbol{\Psi}$ can be sampled indirectly by imposing independent, entry-wise $N(0,1)$ priors on $\mathbf{X}$, corresponding to a uniform prior density on $\mathcal{V}_{K,Q}$. This is a particular case of the parameter expansion strategy used by Jauch et al., (2021); Park, (2025); Meng and Bouchard, (2024).

### v4

# Fast Bayesian Functional Principal Components Analysis

@@ Abstract
Functional Principal Components Analysis (FPCA) is a widely used analytic tool for dimension reduction of functional data. Traditional implementations of FPCA estimate the principal components from the data, then treat these estimates as fixed in subsequent analyses. To account for the uncertainty of PC estimates, we propose FAST, a fully-Bayesian FPCA with three core components: (1) projection of eigenfunctions onto an orthonormal spline basis; (2) efficient sampling of the orthonormal spline coefficient matrix using a parameter expansion scheme based on polar decomposition; and (3) ordering eigenvalues during sampling. Extensive simulation studies show that FAST is very stable and performs better compared to existing methods. FAST is motivated by and applied to a study of the variability in mealtime glucose from the Dietary Approaches to Stop Hypertension for Diabetes Continuous Glucose Monitoring (DASH4D CGM) study. All relevant STAN code and simulation routines are available as supplementary material.
@@ 1 Introduction
Functional principal component analysis (FPCA) is a common analytic approach for high-dimensional applications. FPCA approximates the covariance between all possible pairs of observations by identifying a small set of orthonormal principal components (FPCs). Using FPCs to provide a reduced-rank representation of functional data, rather than fixed (spline) bases (Shi et al.,, 1996; Rice and Wu,, 2001), provides a data-driven representation based on the leading modes of variability (James et al.,, 2000). While the means of estimating these FPCs can vary, e.g. polynomial smoothing (Yao et al.,, 2005), kernel smoothing (Staniswalis et al.,, 1998), or sandwich covariance smoothing (Xiao et al.,, 2016), the key idea remains consistent. Conditioning on the FPCs, functional models become mixed effects models with a small number of random effects and inference can be conducted using existing software (Reiss and Ogden,, 2007; Liu and Guo,, 2012; Di et al.,, 2009; Zhou et al.,, 2010; Shou et al.,, 2015). Frequentist implementations of FPCA follow a two-step approach: (1) diagonalizing a smooth estimator of the observed functional covariance; and then (2) conducting inference conditional on the FPC estimates (Xiao et al.,, 2016; Crainiceanu and Goldsmith,, 2010; Peng and Paul,, 2009). This conditional approach may lead to invalid inference in smaller samples and/or when eigenfunctions explain smaller variance (Goldsmith et al.,, 2013). FPC forms may be falsely identified, or FPC scores with substantial uncertainty may be used in downstream regression.
@@
There are several Bayesian FPCA implementations that model jointly the FPCs and their scores. An early fully-Bayesian FPCA was developed by Goldsmith et al., (2015), who modeled both the fixed and random effects functions using unconstrained splines and applied post-processing to obtain orthonormal FPC samples. Jauch et al., (2021) directly sampled the FPC matrix using a parameter expansion scheme based on polar decomposition. Suarez and Ghosal, (2017) proposed an empirical Bayes approach based on spline expansion which models the coefficient covariance structure. A recent approach used variational inference based on message passing (Nolan et al.,, 2023). Boland et al., (2023) implemented Bayesian FPCA using latent factor methods, also providing useful visualization tools for posterior FPC uncertainty. In the context of sparse functional data, Ye, (2024) projected the FPCs onto an orthonormal spline basis and used Gibbs sampling approach of Hoff, (2009) to sample the spline weights such that FPC orthonormality was preserved. Recently Zhang et al., (2025) proposed a robust implementation based on sequential Monte Carlo designed to handle outliers and sparse observations.
@@
In this paper, we introduce Fast Bayesian FPCA, or FAST for short, which fuses ideas introduced Ye, (2024) and Jauch et al., (2021) and expands them to create a self-contained methodology that is computationally stable, scalable, and extensible to a variety of other data structures and models. FAST leverages orthogonal spline expansion to reduce the dimension of sampled FPC parameters to the corresponding weights, which are then efficiently sampled using a parameter expansion scheme based on polar decomposition. The dimensionality reduction of the posterior space by spline expansion combined with stabilizing order constraints of FPCA ensures that the approach has excellent convergence and mixing properties while being computationally efficient. All core components of FAST are easy to implement in Bayesian software such as STAN (Carpenter et al.,, 2017) and can be extended to include covariates, multilevel, sparse, and multivariate data.
@@ 2.1 FPCA model
The functional data structure has the form $Y_{i}(t),i=1,\ldots,N$ for $t\in\{t_{1},\ldots,t_{M}\}\in[0,1]$, which can be thought of as functions observed on a finite grid. Assuming that $Y_{i}(t)$ are generated by the same Gaussian Process and truncating the decomposition (Karhunen,, 1947; Kosambi,, 1943; Loève,, 1978) produces the standard FPCA model:
@@
Model (1) is a first line approach in the analysis of high dimensional data (Crainiceanu et al.,, 2024) because in many applications $K\ll M$ and model (1) provides a practical, low-dimensional approximation to the original high-dimensional data. Moreover, conditional on the FPCs, $\phi_{k}(\cdot)$, the model becomes a mixed effects model with uncorrelated random effects. In this paper we treat the eigenfunctions $\phi_{k}(t)$ as parameters, estimate them from the data, and account for their uncertainty. FAST provides a stable approach to simulating the posterior distribution of all parameters, including the eigenfunctions.
@@ 2.2 Orthonormal spline expansion
All functional components are expanded using a $Q$-dimensional orthonormal spline basis, such that $Q\geq K$. Orthonormality is defined with respect to the scalar product on $L^{2}([0,1])$, i.e. $\langle f,g\rangle=\int_{0}^{1}f(x)g(x)dx$. Denote by $\mathbf{B}(t)=\{B_{1}(t),\ldots,B_{Q}(t)\}$ the spline basis and by $\mathbf{B}\in\mathbb{R}^{M\times Q}$ the matrix representation of the spline evaluated over an $M$-dimensional grid. Each column of $\mathbf{B}$ corresponds to a spline basis and each row corresponds to a sampling point. We model $\mu(t)=\mathbf{B}(t)w_{\mu}$ and $\phi_{k}(t)=\mathbf{B}(t)\psi_{k}$, where $w_{\mu}$ and $\psi_{k}$ are $Q$-dimensional vectors of spline coefficients. In the data space, these become $\{\mu(t_{1}),\ldots,\mu(t_{M})\}^{t}=\mathbf{B}w_{\mu}$ and $\{\phi_{k}(t_{1}),\ldots,\phi_{k}(t_{M})\}^{t}=\mathbf{B}\psi_{k}$.
@@
We will show that the $\phi_{k}(t)$ functions are orthonormal if and only if $\boldsymbol{\Psi}^{t}\boldsymbol{\Psi}=I_{K\times K}$, where $\mathbf{\Psi}=[\psi_{1}|\ldots,|\psi_{K}]$ is the $Q\times K$ dimensional matrix obtained by column binding the $\psi_{k}$. Indeed, as $\mathbf{B}(t)$ is an orthonormal spline basis, $\langle B_{i}(t),B_{j}(t)\rangle=1$ if $i=j$ and 0 otherwise. Therefore, $\langle\phi_{k}(t),\phi_{k^{\prime}}(t)\rangle$ is equal to
$\int_{0}^{1}[\mathbf{B}(t)\psi_{k}][\mathbf{B}(t)\psi_{k^{\prime}}]dt=\sum_{i=1}^{Q}\sum_{j=1}^{Q}\psi_{k,i}\psi_{k^{\prime},j}\bigl[\int_{0}^{1}B_{i}(t)B_{j}(t)dt\bigr]=\psi_{k}^{t}\psi_{k^{\prime}}\;.$
@@
The quantity $\psi_{k}^{t}\psi_{k^{\prime}}=1$ when $k=k^{\prime}$ and 0 otherwise if and only if $\boldsymbol{\Psi}$ is orthonormal. Therefore, the eigenfunctions $\phi_{k}(t)$ are orthonormal in $L^{2}([0,1])$ if and only if the low-dimensional spline coefficients are orthonormal in $\mathbb{R}^{Q}$. This is a crucial point of our methodology, as the orthonormality of the eigenfunctions (in infinite or high dimensional spaces) is controlled by orthonormality of vectors in a small dimensional space. Because it is much easier to sample from a small dimensional space, this projection is one of the main reasons the FAST method is stable and computationally efficient.
@@
In functional data analysis, orthonormality in the theoretical $L^{2}([0,1])$ functional space is often confounded with orthonormality in the $\mathbb{R}^{M}$ observed vector space. For a fixed $M$ and equally-spaced observations, the inner product on $\mathbb{R}^{M}$ is the Riemann sum approximation (up to the scaling factor $1/M$) of the functional inner product. However, this approximation depends on the density of the observation in the functional domain and needs to be re-adjusted when the number of sampling points is changed. FAST can be used in the vector space as well by choosing the matrix $\mathbf{B}$ such that $\mathbf{B}^{t}\mathbf{B}=\mathbf{I}_{Q}$. The resulting FPC matrix $\mathbf{B}\boldsymbol{\Psi}$ is then a product of orthonormal matrices. To maintain robustness to different $M$ and sampling regimes, we focus here on functional orthonormality. However, our software implementation can use vector orthormality, as well.
@@
The basis dimension $Q$ must be sufficiently large to capture the maximum complexity of the subject-specific functions. We follow here the standard recommendations for penalized splines and use $Q\in[20,40]$ (Ruppert,, 2002), which performed well in our tests (see Supplement Sections S5.2). The choice of the basis matrix $\mathbf{B}$ can substantially impact the computational stability and efficiency of FAST. While, in theory, any spline matrix $\mathbf{B}$ which is orthonormal in $L^{2}[0,1]$ could be used, in practice some approaches have better computational properties and interpretability. We choose as our basis ‘Splinets’, an orthonormalized B-spline basis introduced by Liu et al., (2020) to maintain temporal localization. We augment their basis with a slope and intercept, appropriately orthonormalized.
@@ 2.3 Priors
We add smoothing penalties on $\mu(t)$ and the FPCs $\phi_{k}(t)$ to the target posterior, which have form $\alpha\int f^{2}(t)dt+(1-\alpha)\int\{f^{\prime\prime}(t)\}^{2}dt$ for generic function $f(\cdot)$. This penalty, used by Goldsmith et al., (2015), contains the fixed weighting parameter $\alpha$, where $\alpha=0$ corresponds to the famous penalized spline prior (Craven and Wahba,, 1979; O’Sullivan,, 1986). If $f(t)=\mathbf{B}(t)\theta$, then the penalty is a quadratic form of $\theta$ with associated matrix $\mathbf{P}_{\alpha}$. Supplement Section S1 covers the derivation of $\mathbf{P}_{\alpha}$. We use $\alpha=0.1$ for consistency with existing implementations, but FAST can handle any $\alpha\in[0,1]$. Supplement Section S5.3 indicates robustness to reasonable choice of $\alpha$. The quadratic penalties are equivalent to Normal priors when $\mathbf{P}_{\alpha}$ is non-degenerate (Brumback et al.,, 1999; Ruppert et al.,, 2003; Wood,, 2006). The penalties $g_{\mu}(w_{\mu}),g_{\phi_{k}}(\psi_{k})$ have the following forms:
$g_{\mu}(w_{\mu})=h_{\mu}^{\text{R}(\mathbf{P}_{\alpha})/2}\exp\left\{-\frac{h_{\mu}}{2}w^{t}_{\mu}\mathbf{P}_{\alpha}w_{\mu}\right\}\;,\;\;g_{\phi_{k}}(\psi_{k})=h_{k}^{\text{R}(\mathbf{P}_{\alpha})/2}\exp\left\{-\frac{h_{k}}{2}\psi^{t}_{k}\mathbf{P}_{\alpha}\psi_{k}\right\}\;,$
where $\text{R}(\mathbf{P}_{\alpha})$ is the rank of the penalty $\mathbf{P}_{\alpha}$ and $h_{\mu},h_{k}>0$ are the smoothing parameters for $\mu(\cdot)$ and $\phi_{(}\cdot)$, respectively. FAST uses separate parameters for each functional component, in-line with Ye, (2024) and Goldsmith et al., (2015).
We assume that the inverse variance components, including the mean smoothing parameter, have Gamma priors with the shape and rate parameters equal to $0.001$. For the FPC smoothing parameters, we choose shape parameters of $0.01$ and rate parameters of equal to the trace of $\mathbf{P}_{\alpha}/2$ plus $0.01$ to ensure a proper prior (see Result 1). For discussion on the choice of Gamma priors for inverse variance components, see Crainiceanu et al., (2005); Crainiceanu and Goldsmith, (2010).
Consider the joint prior $g(\boldsymbol{\Psi},\mathbf{H})=g_{\psi}(\boldsymbol{\Psi}|\mathbf{H})g_{h}(\mathbf{H})$ on the FPCs ($\Psi$) and the smoothing parameters $\mathbf{H}=\operatorname{diag}(h_{1},\ldots,h_{K})$. Here $g_{\psi}(\cdot)$ is the conditional prior of $\boldsymbol{\Psi}$ given the smoothing parameter matrix $\mathbf{H}$ and $g_{h}(\cdot)$ is the product of independent Gamma priors on $h_{k}$. We define the conditional prior $g_{\psi}(\boldsymbol{\Psi}|\mathbf{H})$ as follows:
$g_{\psi}(\boldsymbol{\Psi}|\mathbf{H})\propto\frac{\prod_{i=1}^{K}h_{i}^{{\rm R}(\mathbf{P}_{\alpha})/2}}{{\rm Vol}(\mathcal{V}_{K,Q})}\times\text{\rm exp}\{-{\rm trace}(\mathbf{H\boldsymbol{\Psi}^{t}\mathbf{P}_{\alpha}}\boldsymbol{\Psi}/2)\}\;,$
where ${\rm Vol}(\mathcal{V}_{K,Q})$ is the volume of the Stiefel manifold $\mathcal{V}_{K,Q}$ and “trace($\boldsymbol{A}$)” denotes the trace of the matrix $\boldsymbol{A}$. The conditional density $g_{\psi}(\boldsymbol{\Psi}|\mathbf{H})$ is a combination of a uniform and smoothing prior on individual PCs. Implementing this prior in Bayesian software is straightforward.
Result 1.
The prior distribution $g(\boldsymbol{\Psi},\mathbf{H})=g_{\psi}(\boldsymbol{\Psi}|\mathbf{H})g_{h}(\mathbf{H})$ is proper if $\beta_{\psi}>{\rm trace}(\mathbf{P}_{\alpha})/2$, where $\beta_{\psi}$ is the rate parameter for the Gamma priors on the smoothing parameters $h_{k}$.
Result 1 provides sufficient conditions under which the joint prior $g(\boldsymbol{\Psi},\mathbf{H})$ is proper; proof can be found in Supplement Section S2.
@@ 2.4 Polar decomposition
The conditional posterior distributions for all parameters except $\boldsymbol{\Psi}$ have analytical forms; see Supplement Section S3. The $Q\times K$ dimensional matrix parameter $\boldsymbol{\Psi}$ is orthonormal and, hence, belongs to the Stiefel manifold (James,, 1976), which has a finite volume and is denoted as $\mathcal{V}_{K,Q}$ (Chikuse,, 2003; Mardia and Khatri,, 1977). Our prior on $\mathbf{\Psi}$, conditional on the smoothing parameters $\{h_{1},\ldots,h_{K}\}$, is a prior on $\mathcal{V}_{K,Q}$.
Result 2.
@@
The distribution in Result 2 does not follow a known form, though it is close to the generalized Langevin-Bingham family (Hoff,, 2009). We sample this parameter indirectly using a parameter expansion based on the polar decomposition. The polar decomposition of the $Q\times K$ dimensional matrix $\mathbf{X}$ is $\mathbf{X}=\mathbf{UP}$, where $\mathbf{U}\in\mathbb{R}^{Q\times K}$ is orthonormal (on the Stiefel manifold $\mathcal{V}_{K,Q}$) and $\mathbf{P}\in\mathbb{R}^{K\times K}$ is positive semi-definite. The polar decomposition is unique when $\mathbf{X}$ is full rank $K$. Crucially, if $\mathbf{X}$ has independent $N(0,1)$ entries, then the matrix $\mathbf{U}$ is uniformly distributed over the Stiefel manifold $\mathcal{V}_{K,Q}$ (Chikuse,, 2003). Therefore, the uniform component of the prior on $\boldsymbol{\Psi}$ can be incorporated by imposing independent, entry-wise $N(0,1)$ priors on $\mathbf{X}$. This is a particular case of the parameter expansion strategy used by Jauch et al., (2021); Park, (2025); Meng and Bouchard, (2024).